If the complement of an angle is one-tenth of the supplement of the angle, find the measure of the angle.
Answer by dabanfield(803)

If the complement of an angle is one-tenth of the supplement of the angle, find the measure of the angle.
Let x be the unknown angle. Then x's complement is 90-x and its supplement is 180 - x. So we have:
90-x = (1/10)*(180-x)
90-x = 18 - .1*x
Solve the above for x.